# Fact or Fiction: Can Truthful Mechanisms Eliminate Federated Free Riding?

### v1

@@ Abstract
Standard federated learning (FL) approaches are vulnerable to the free-rider dilemma: participating agents can contribute little to nothing yet receive a well-trained aggregated model. While prior mechanisms attempt to solve the free-rider dilemma, none have addressed the issue of truthfulness. In practice, adversarial agents can provide false information to the server in order to cheat its way out of contributing to federated training. In an effort to make free-riding-averse federated mechanisms truthful, and consequently less prone to breaking down in practice, we propose Fact. Fact is the first federated mechanism that: (1) eliminates federated free riding by using a penalty system, (2) ensures agents provide truthful information by creating a competitive environment, and (3) encourages agent participation by offering better performance than training alone. Empirically, Fact avoids free-riding when agents are untruthful, and reduces agent loss by over $4$x. Code for Fact is found on GitHub at https://github.com/marcobornstein/FACT.
@@ 6 Experimental Results
Fact Enforces Truthfulness. The results of Figure 3 back the theoretical result of Theorem 1, which states that agents maximize their expected reward (i.e., net reduction in loss versus local training) when truthfully reporting their cost. Across both homogeneous (iid) and heterogeneous (n-iid) agent data distributions, agents’ net improvement in loss peaks when reporting their true cost (0% added) and monotonically decreases as agents report an inflated or deflated true cost. Thus, the sandwich-style truthfulness competition Fact employs disincentivizes agents from inflating or deflating their true costs. This is important, as agents are no longer incentivized to lie to the server about their true cost in order to free-ride without the free-rider penalty in Equation (4) being imposed. Net improvement in loss grows as data distributions become increasingly non-iid. FL is robust enough to deal with the increasingly non-iid datasets while local training suffers (Figures 4 & 5). Therefore, the gap between FL and local training grows, and thus so too does the reward in Fact.
Fact Reduces Agent Loss. Agents participating in Fact on average achieve a reduction in total loss compared to training on their own when using the same amount of data. Consequently, participating in Fact is individually rational: agents at worst receive loss equivalent to local training and on average receive improved loss (Figure 2). Traditional FL achieves a lower average agent loss than Fact when all agents use as much data as they would locally. However, as shown in Theorem 3, this is unrealistic. Traditional FL emits a catastrophic free-riding optimum. The results of traditional FL in Figure 2 is the upper bound on achievable loss by agents, only attainable if agents refused to free ride. The gap in loss between Fact and the upper bound of traditional FL is the cost of agent untruthfulness and their inclination to free ride. The gap is the direct result of the penalty terms on free riding $P_{fr}$ and truthfulness $P_{ct}$ found in Equation (11). Average local loss increases as data distributions become increasingly non-iid, while traditional FL remains robust to the distribution shift (Figures 4 & 5). Since the agents in Fact who lose the truthfulness competition receive a loss equivalent to local training, the loss of Fact also rises in proportion to the local training loss.
@@ Appendix B Additional Experimental Results
Test Loss Plots. We include the test loss plots, including error bars, below for federated and local training for both CIFAR-10 and MNIST. We ran each of our experiments three times. As expected, federated training outperforms local training in the iid (left), mild non-iid (middle) and strong non-iid (right) settings. Federated training, via FedAvg, is more robust to non-iid distributions than local training. As detailed in Section 6, our non-iid distributions are Dirichlet with parameters $\alpha=0.3,0.6$.
Figure 4: Loss Plots for CIFAR10.
Figure 5: Loss Plots for MNIST.
@@ Appendix C Proofs
$\frac{\partial}{\partial c}\bigg{[}\underbrace{\frac{\gamma\sigma^{2}L}{2m_{i}^{*}}+c_{i}m_{i}^{*}}_{\ell_{i,l}(m_{i}^{*})}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}$ $=0,$
$-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}\upsilon(c)\bigg{]}$ $=0,$
@@
Appendix D Impact Statement

### v2

@@ Abstract
Standard federated learning (FL) approaches are vulnerable to the free-rider dilemma: participating agents can contribute little to nothing yet receive a well-trained aggregated model. While prior mechanisms attempt to solve the free-rider dilemma, none have addressed the issue of truthfulness. In practice, adversarial agents can provide false information to the server in order to cheat its way out of contributing to federated training. In an effort to make free-riding-averse federated mechanisms truthful, and consequently less prone to breaking down in practice, we propose Fact. Fact is the first federated mechanism that: (1) eliminates federated free riding by using a penalty system, (2) ensures agents provide truthful information by creating a competitive environment, and (3) encourages agent participation by offering better performance than training alone. Empirically, Fact avoids free-riding when agents are untruthful, and reduces agent loss by over $4$x.
@@ 6 Experimental Results
Fact Enforces Truthfulness. The results of Figure 3 back the theoretical result of Theorem 1, which states that agents maximize their expected reward (i.e., net reduction in loss versus local training) when truthfully reporting their cost. Across both homogeneous (iid) and heterogeneous (n-iid) agent data distributions, agents’ net improvement in loss peaks when reporting their true cost (0% added) and monotonically decreases as agents report an inflated or deflated true cost. Thus, the sandwich-style truthfulness competition Fact employs disincentivizes agents from inflating or deflating their true costs. This is important, as agents are no longer incentivized to lie to the server about their true cost in order to free-ride without the free-rider penalty in Equation (4) being imposed. Net improvement in loss grows as data distributions become increasingly non-iid. FL is robust enough to deal with the increasingly non-iid datasets while local training suffers (Figures 5 & 6). Therefore, the gap between FL and local training grows, and thus so too does the reward in Fact.
Fact Reduces Agent Loss. Agents participating in Fact on average achieve a reduction in total loss compared to training on their own when using the same amount of data. Consequently, participating in Fact is individually rational: agents at worst receive loss equivalent to local training and on average receive improved loss (Figure 2). Traditional FL achieves a lower average agent loss than Fact when all agents use as much data as they would locally. However, as shown in Theorem 3, this is unrealistic. Traditional FL emits a catastrophic free-riding optimum. The results of traditional FL in Figure 2 is the upper bound on achievable loss by agents, only attainable if agents refused to free ride. The gap in loss between Fact and the upper bound of traditional FL is the cost of agent untruthfulness and their inclination to free ride. The gap is the direct result of the penalty terms on free riding $P_{fr}$ and truthfulness $P_{ct}$ found in Equation (11). Average local loss increases as data distributions become increasingly non-iid, while traditional FL remains robust to the distribution shift (Figures 5 & 6). Since the agents in Fact who lose the truthfulness competition receive a loss equivalent to local training, the loss of Fact also rises in proportion to the local training loss.
@@
6.1 Real-World Case Study: Inter-Hospital Skin Cancer Diagnosis
Figure 4: Fact Eliminates Free Riding in Realistic Settings. When training an image classifier for diagnosing skin cancer, agents participating in FACT achieve much lower loss ($66\%$ less) than if they did not participate (left). Agents maximize their improvement in loss over local training when they are truthful; reporting inflated or deflated costs diminishes improvement in loss (middle). Agents minimize penalties when using their locally optimal amount of data ($m^{*}=801$) for training (right).
We consider a realistic situation where a consortium of hospitals seek collaboration to train a model, privately in a FL manner, that can diagnose skin cancer. One of the hospitals is smaller and resource-constrained. It is difficult, but not impossible, for this hospital to collect more data for training. Thus, in the absence of a truthful FL mechanism, the smaller hospital could over-report its collection costs to the server in order to contribute little to no data towards training while still reaping the rewards of a well-trained global model.
To test how Fact deals with this scenario, we train an image classification model on the HAM10000 [31] dataset. HAM10000 consists of a collection of dermatoscopic images of pigmented lesions. The goal is to train a model which can perform automated diagnosis of pigmented skin lesions, including melanoma (skin cancer). Our setup is similar to the experimental setup detailed in the previous section. One difference is that we fine-tune a ResNet50 model on HAM10000 that is pre-trained on ImageNet (a realistic approach for a hospital). HAM10000 is an imbalanced dataset, and evenly partitioning $80\%$ of the data amongst 10 devices, as their local training sets, further exacerbates the non-iidness of the data. This too is realistic, as data is often non-iid amongst FL agents. We use the Adam optimizer with a learning rate of 1e-3 and batch size of 128.
As expected, Fact provably dissuades any hospital from lying about its cost in order to reduce their data contribution level. Agents are incentivized to be truthful; reduction in loss is maximized when agents truthfully report their costs (see the middle plot in Figure 4). As a result, a smaller, resourced-constrained hospital would still contribute its locally optimal amount of data, or face harsh penalties (see the right plot in Figure 4). Finally, all participating hospitals are still better off participating in Fact than training alone. Participating agent loss is reduced by nearly $66\%$ compared to local training (see the left plot in Figure 4). The effectiveness of Fact to dissuade free-riding in the real-world setting of training a skin cancer classifier across hospitals indicates that it will be successful in other real-world settings.
@@ Appendix B Additional Experimental Results
Test Loss and Accuracy Plots. We include the test loss and accuracy plots, including error bars, below for federated and local training for CIFAR-10, MNIST, and HAM10000. We run each experiment three times. As expected, FL outperforms local training in the iid (left), mild non-iid (middle) and strong non-iid (right) settings. Federated training, via FedAvg, is more robust to non-iid distributions than local training. As detailed in Section 6, our non-iid distributions are Dirichlet with $\alpha=0.3,0.6$.
Figure 5: Test Loss for CIFAR10 (top) and MNIST (bottom) in Heterogeneous Settings. FL outperforms local training on iid (left) and mild (middle) & strong (right) non-iid Dirichlet settings.
Figure 6: Test Accuracy for CIFAR10 (top) and MNIST (bottom) in Heterogeneous Settings. FL outperforms local training on iid (left) and mild (middle) & strong (right) non-iid Dirichlet settings.
Figure 7: HAM10000 Test Loss (Left) and Accuracy (Right) for Federated and Local Training.
@@ Appendix C Proofs
$\frac{\partial}{\partial c}\bigg{[}\underbrace{\frac{\gamma\sigma^{2}L}{2m_{i}^{*}}+c_{i}m_{i}^{*}}_{\ell_{i,l}(m_{i}^{*})}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}$ $=0\longrightarrow-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}\upsilon(c)\bigg{]}=0,$
@@
Appendix D Impact Statement & Limitations
@@
Our work faces limitations when agents and the central server act maliciously. On the agent side, we assume that agents are rational and do not collude with one another. Our future research direction is to prove that equilibria for Fact exist when agents are boundedly rational or colluding. On the server side, we assume that the central server acts honestly. In settings where the server cannot be trusted, new incentives or avenues must be built in order to ensure server honesty.